Make New Entry,
Make Followup Entry
User name brads
Log entry time 21:56:14 on March 12, 2012
Entry number 366045
This entry is a followup to: 365730
keyword=GEp/G2p status update -- Monday
GEp production running for their first point is complete. Overall,
data taking went very smoothly. Production running ended around noon
took beam away for a spot move. Hall A went to Controlled Access and stayed
there for a TE measurement. The TE took longer than anticipated so we're
running a few hours later than we had initially planned, but we will still be
able to achieve our goals for SWING and OWL.
A run-plan for tonight, Wed, and Thursday is posted in
366034. As I write this (9:30pm) we are wrapping up the BCM calibration.
We will continue with a Moller, follow up with BPM calibrations (HARP 1HA05
was repaired during today's access), and a BCM calorimeter run. The
OWL shift will be spent taking inelastic data.
The target's slow depolarization mystery is explained by a BCM calibration
over estimating our beam current on target. It was closer to 15--20nA, not
the 55nA reported by 1H00. This was not a problem for us since we
DAQ rate limited -- in fact, the persistent high polarization really improved
the figure of merit for this GEp setting.
We were unable to run straight through optics in this chicane
initially planned. We shuffled the runplan and will be using the Tuesday
chicane move to go to the Straight Through setting. That time overlaps with
an extended Beam Study period, so we don't lose any beam time. SWING and OWL
will be spent on our planned Optics studies and possible dilution/ background
We will make a second chicane move on Wednesday DAY to Position #5 (2.2 GeV,
Target field = 2.5 T, Target orientation = 90 deg). This will set us up
for our first g2p production running.
Issues to watch for (these are also on the whiteboard):
- Shift staffing is a still a problem! It's looking better in the short
term, but there are many vacancies starting March 18 -- less than a week
Come on folks, step up and sign up! Really -- no kidding.
- Beam leakage from Hall C is well under control now. We need to be
vigilant and make sure that Hall A Laser slit is set to 14 for our production
running. The shift crew are requested follow a quick bleedthrough check
every 4--5 hours just to be safe.
- The IHWP will be flipped at Hall C's request every 4 hours.
Hall C and/or MCC should inform us when this happens so we
can log the change and stop/start our runs. Hall C would like to shorten
this flip interval in the future -- we are working on a procedure and
frequency that will work for both Halls.
- The target magnet current spontaneously drops by ~0.5A occasionally
(look in Halog for "Target Current" entries). The Target operator needs
to watch for these and know how to correct the issue -- ask the TO how to
do this when you come on shift if you don't already know. This is likely to
be a bigger problem when we return to higher currents. (We will attempt to
add some shielding around relevant electronics during Tuesday's access.)
A copy of this log entry has been emailed to: g2p